OpenStudy (anonymous):

What translation has taken place in the graph if f(x) = x^3-4 from the graph of the parent function?

OpenStudy (zzr0ck3r):

g(x) = f(x)+b takes f(x)(the parent function) and moves it up b units

OpenStudy (zzr0ck3r):

if b is negative is movies it down b units

OpenStudy (itsonlycdeee):

The parent function is f(x) = x^3, since the one in this case is f(x)=x^3-4, 4 determines the direction of the vertical shift. @zzr0ck3r thanks aha